Fermi National Laboratory


Accelerator Update for September 2, 2005 - September 5, 2005

Friday, September 2
The day shift began with Operations monitoring store 4363, the Recycler Ring (RR) stash at 48.0 E10, the Antiproton Source (Pbar) stack at 44.8 E10, and with MiniBooNE, NuMI, and the Meson MCenter down, due to problems with the 750Kev buncher.

Beam resumes at 8:19 AM.

End of store studies in the Tevatron (Tev) begin at 10:06 AM.

Store 4363 was terminated at 12:01 PM.

At 12:13 PM, CDF and D0 are ready to access their collision halls.

Beam was turned off at 1:04 PM to change the Power Amp on Linac RF station #1 (LRF1).

CDF and D0 accesses were completed at 2:00 PM.

At 3:30 PM the Linac RF station was repaired, stacking resumes, along with beam to NuMI, MiniBooNE, and MCenter.

Tevatron studies begin at 5:25 PM.

Shot setup begins at 9:51 PM.

Saturday, September 3
The midnight shift began with Operations in shot setup, and with MiniBooNE, NuMI taking beam.

Operations established store 4367 at 12:29 AM with an initial luminosity of 95.2 E30.

Pbar returns to stacking at 12:30 AM, beam to MCenter.

Operations put the Recycler into shot setup at 3:51 AM to stash antiprotons.

Lost the Main Injector (MI) power supply loop at 4:40 AM, fortunately Pbars were not being transferred at the time. We were unable to reset the loop and power supply experts were called in. MiniBooNE continues to take beam, NuMI, Mcenter, and Pbar production are off.

The Main Injector power supply loop was back on by 10:57 AM.

The Tev store was lost at 11:02 AM, due to a non-existant vacuum problem at B3. The status chassis indicated a vacuum valve had closed, even though no valves were found closed.

At 11:34 AM we had a vacuum burst at F0 that cascaded through the P150, A150, and NuMI lines.

Beam returns to NuMI and MCenter at 12:15 PM.

Pbar returns to stacking at 12:36 PM.

Since the Tev was off, the run Coordinator is allowing Tev access for necessary repairs.

At 2:43 PM a 2 inch cap had blown off one of the pond water cooling lines at C0. This was causing water to spray all over one of the operating compressors. The compressor was turned off and the cap was replaced. The compressor was back on by 5:00 PM.

At 2:57 PM, MKS02 tripped off causing MCenter to be turned off. Experts were called to facilitate repairs.

All Tev accesses were completed by 5:10 PM.

MKS02 (a PFN cable had to be reterminated) was back on by 5:39 PM. Beam returns to MCenter.

At 10:19 PM all Tev studies were complete.

shot setup begins at 11:33 PM.

Sunday, September 4
The midnight shift began with Operations in shot setup, and with MiniBooNE, NuMI taking beam.

Operations established store 4369 at 2:44 AM with an initial luminosity of 80.4 E30.

Operations put the Recycler into shot setup at 6:58 AM to stash antiprotons.

Pbar returns to stacking at 7:47 AM, beam also returns to Mcenter.

At 10:57 AM, a quad power supply D:Q801 started to regulate 80 amps higher than normal, thus causing the Pbar production rate to drop. An expert was called in, the logic card was replaced, and normal stacking resumed at 12:10 PM.

Operations put the Recycler into shot setup at 1:00 PM to stash antiprotons.

Pbar returns to stacking at 2:00 PM, beam also returns to MCenter.

Monday, September 5
The midnight shift began with Operations monitoring store 4369, the Recycler Ring (RR) stash at 99.5 E10, Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Store 4369 was terminated at 2:07AM.

Shot setup begins at 2:19AM.

Operations established store 4370 at 5:18 AM with an initial luminosity of 114.12 E30.

Operations put the Recycler into shot setup at 6:06 AM to stash antiprotons.

Pbar returns to stacking at 6:54 AM, beam returns to MCenter.

The Plan for Monday
Keep the store, continue stacking and shooting to the Recycler, run beam to NuMI, MiniBoone, and MCenter.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/6/2005   email Fermilab